# The Map Equation Goes Neural: Mapping Network Flows with Graph Neural Networks

### v3

@@ B.2 Tabulated AMI results on real-world networks
Method Arch. CS Physics Cora ML Wiki CS ogb-arxiv
Neuromap LIN $\underline{76.7}\pm 1.4$ $56.5\pm 1.2$ $8.0\pm 4.9$ $41.2\pm 2.0$ $9.9\pm 9.1$
MLP $\textbf{78.1}\pm 2.0$ $55.9\pm 2.2$ $31.0\pm 2.6$ $\textbf{44.8}\pm 2.8$ $32.9\pm 1.3$
GCN $70.4\pm 1.6$ $50.3\pm 1.4$ $36.8\pm 1.9$ $42.6\pm 3.0$ $42.1\pm 1.4$
GIN $69.1\pm 2.8$ $52.7\pm 2.7$ $35.7\pm 1.7$ $35.6\pm 5.0$ $26.7\pm 3.3$
SAGE $70.9\pm 2.4$ $51.9\pm 2.2$ $38.1\pm 1.6$ $43.3\pm 2.9$ $\textbf{43.9}\pm 0.6$
DMoN LIN $30.4\pm 6.7$ $15.4\pm 9.6$ $1.6\pm 2.1$ $6.8\pm 4.3$ $4.9\pm 3.6$
MLP $50.9\pm 0.9$ $32.5\pm 0.7$ $10.1\pm 1.3$ $28.6\pm 1.2$ $18.9\pm 0.8$
GCN $52.2\pm 0.4$ $31.5\pm 0.4$ $29.8\pm 1.1$ $34.7\pm 0.4$ $35.4\pm 0.3$
GIN $50.0\pm 0.7$ $30.3\pm 0.8$ $32.0\pm 1.2$ $30.0\pm 1.1$ $23.7\pm 0.9$
SAGE $52.1\pm 0.5$ $31.9\pm 0.4$ $29.5\pm 1.4$ $34.7\pm 0.3$ $33.4\pm 0.7$
NOCD LIN $40.2\pm 9.8$ $19.3\pm 7.0$ $3.3\pm 1.1$ $10.8\pm 3.1$ $4.2\pm 2.5$
MLP $73.9\pm 1.7$ $50.2\pm 1.0$ $\textbf{46.5}\pm 1.4$ $40.9\pm 1.2$ $9.7\pm 2.3$
GCN $63.3\pm 1.0$ $42.0\pm 0.7$ $39.1\pm 1.0$ $\underline{44.2}\pm 0.6$ $\underline{43.8}\pm 0.4$
GIN $63.7\pm 0.9$ $48.3\pm 1.5$ $\underline{43.3}\pm 1.2$ $37.6\pm 1.1$ OOM
SAGE $63.7\pm 0.8$ $42.4\pm 0.7$ $40.6\pm 1.1$ $42.4\pm 0.4$ $43.1\pm 0.3$
DiffPool LIN $22.2\pm 14.3$ $18.9\pm 15.9$ $0.0\pm 0.1$ $2.4\pm 2.1$ OOM
MLP $64.5\pm 4.8$ $49.1\pm 4.4$ $9.5\pm 1.2$ $25.8\pm 2.6$ OOM
GCN $73.3\pm 2.0$ $\textbf{62.4}\pm 3.6$ $33.5\pm 1.2$ $35.7\pm 3.8$ OOM
GIN $50.3\pm 3.9$ $41.0\pm 3.9$ $31.6\pm 2.8$ $11.3\pm 3.1$ OOM
SAGE $71.4\pm 3.4$ $\underline{59.2}\pm 3.3$ $29.6\pm 2.2$ $38.1\pm 2.8$ OOM
MinCut LIN $57.5\pm 0.6$ $36.8\pm 0.9$ $30.4\pm 1.2$ $16.0\pm 1.4$ $3.8\pm 3.5$
MLP $53.9\pm 6.4$ $29.4\pm 9.9$ $39.6\pm 1.3$ $28.1\pm 4.2$ $8.0\pm 2.7$
GCN $44.7\pm 3.3$ $32.2\pm 6.8$ $26.2\pm 0.7$ $32.1\pm 0.5$ $33.9\pm 4.7$
GIN $32.7\pm 19.5$ $51.7\pm 5.4$ $6.9\pm 4.5$ $1.3\pm 6.2$ $10.3\pm 7.3$
SAGE $47.1\pm 0.5$ $30.5\pm 2.4$ $35.9\pm 0.7$ $32.3\pm 0.5$ $33.0\pm 1.5$
Ortho LIN $45.4\pm 2.0$ $26.7\pm 1.5$ $15.6\pm 1.1$ $29.6\pm 0.9$ $19.1\pm 0.5$
MLP $56.3\pm 1.7$ $32.0\pm 1.1$ $5.3\pm 0.5$ $28.8\pm 0.9$ $18.5\pm 0.4$
GCN $51.6\pm 1.0$ $32.4\pm 1.0$ $19.2\pm 1.1$ $32.1\pm 0.5$ $34.3\pm 0.7$
GIN $44.1\pm 0.6$ $25.0\pm 0.7$ $28.7\pm 1.1$ $29.4\pm 0.6$ $24.1\pm 0.8$
SAGE $52.8\pm 1.2$ $34.4\pm 1.1$ $16.3\pm 1.1$ $31.7\pm 0.5$ $24.9\pm 0.3$
Infomap $40.6\pm 0.2$ $26.7\pm 0.1$ $35.7\pm 0.3$ $37.9\pm 0.2$ $35.9\pm 0.1$
B.3 Tabulated number of detected communities on real-world networks

### v4

@@ C.2 Tabulated AMI results on real-world networks
Method Arch. Cora CiteSeer PubMed PC Photo CS Physics Cora ML Wiki CS ogb-arxiv
Neuromap LIN $31.9\pm 7.9$ $19.5\pm 1.6$ $17.2\pm 12.2$ $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $\underline{76.7\pm 1.4}$ $56.5\pm 1.2$ $24.4\pm 11.9$ $41.9\pm 2.3$ $4.7\pm 7.7$
MLP $36.6\pm 2.4$ $15.8\pm 1.1$ $23.2\pm 1.8$ $22.4\pm 8.6$ $39.8\pm 11.2$ $\mathbf{78.1\pm 2.0}$ $55.9\pm 2.2$ $34.5\pm 1.6$ $43.9\pm 2.4$ $32.9\pm 1.3$
GCN $38.5\pm 2.0$ $16.2\pm 1.3$ $20.8\pm 1.2$ $\mathbf{52.1\pm 2.3}$ $\mathbf{65.0\pm 3.8}$ $70.4\pm 1.6$ $50.3\pm 1.4$ $36.4\pm 1.2$ $41.1\pm 3.2$ $42.3\pm 1.7$
GIN $\mathbf{46.1\pm 2.0}$ $23.1\pm 1.2$ $\mathbf{24.6\pm 1.7}$ $43.2\pm 2.4$ $58.2\pm 6.0$ $69.1\pm 2.8$ $52.7\pm 2.7$ $37.5\pm 1.9$ $34.9\pm 4.5$ $28.0\pm 2.8$
SAGE $41.6\pm 1.9$ $18.3\pm 1.2$ $20.4\pm 1.0$ $43.6\pm 4.8$ $48.1\pm 5.8$ $70.9\pm 2.4$ $51.9\pm 2.2$ $39.1\pm 2.1$ $\mathbf{44.8\pm 2.2}$ $\mathbf{43.6\pm 1.0}$
DMoN LIN $1.2\pm 1.6$ $3.2\pm 1.6$ $0.7\pm 1.0$ $3.4\pm 2.9$ $3.7\pm 4.0$ $25.2\pm 10.8$ $5.7\pm 7.3$ $0.4\pm 0.8$ $4.8\pm 3.9$ $4.0\pm 3.6$
MLP $36.6\pm 1.2$ $22.5\pm 0.9$ $6.7\pm 4.1$ $15.4\pm 11.6$ $43.5\pm 8.9$ $40.2\pm 6.5$ $34.7\pm 3.5$ $36.2\pm 1.2$ $31.5\pm 8.0$ $12.8\pm 1.7$
GCN $38.6\pm 0.9$ $20.4\pm 0.9$ $18.3\pm 0.2$ $43.8\pm 0.3$ $50.5\pm 0.3$ $53.1\pm 0.2$ $33.4\pm 0.2$ $37.7\pm 0.9$ $36.7\pm 0.4$ $38.4\pm 0.2$
GIN $37.2\pm 1.1$ $19.5\pm 0.9$ $17.0\pm 0.3$ $35.4\pm 2.3$ $45.6\pm 0.8$ $49.9\pm 0.4$ $29.8\pm 0.4$ $34.7\pm 1.5$ $29.7\pm 0.8$ $26.2\pm 3.2$
SAGE $39.1\pm 1.0$ $21.3\pm 1.3$ $17.7\pm 0.2$ $37.4\pm 11.9$ $50.8\pm 0.3$ $53.1\pm 0.3$ $33.6\pm 0.2$ $40.5\pm 1.0$ $36.4\pm 0.4$ $38.1\pm 0.3$
NOCD LIN $5.1\pm 1.9$ $2.4\pm 1.3$ $3.7\pm 1.7$ $1.2\pm 2.7$ $0.1\pm 0.3$ $40.2\pm 9.8$ $19.3\pm 7.0$ $3.3\pm 1.1$ $10.8\pm 3.1$ $4.2\pm 2.5$
MLP $\mathbf{46.1\pm 1.1}$ $\mathbf{34.8\pm 0.9}$ $20.5\pm 2.4$ $45.5\pm 2.5$ $60.5\pm 3.4$ $73.9\pm 1.7$ $50.2\pm 1.0$ $\mathbf{46.5\pm 1.4}$ $40.9\pm 1.2$ $9.7\pm 2.3$
GCN $41.1\pm 0.9$ $20.2\pm 0.8$ $21.2\pm 0.6$ $\mathbf{51.9\pm 1.9}$ $63.2\pm 0.9$ $63.3\pm 1.0$ $42.0\pm 0.7$ $39.1\pm 1.0$ $\mathbf{44.2\pm 0.6}$ $\mathbf{43.8\pm 0.4}$
GIN $43.4\pm 1.1$ $\underline{26.4\pm 1.2}$ $21.3\pm 0.6$ $34.2\pm 1.9$ $55.3\pm 1.9$ $63.7\pm 0.9$ $48.3\pm 1.5$ $43.3\pm 1.2$ $37.6\pm 1.1$ $28.7\pm 1.2$
SAGE $40.7\pm 0.9$ $19.8\pm 0.8$ $19.6\pm 0.4$ $\mathbf{51.1\pm 1.6}$ $\mathbf{63.7\pm 1.6}$ $63.7\pm 0.8$ $42.4\pm 0.7$ $40.6\pm 1.1$ $42.4\pm 0.4$ $43.1\pm 0.3$
DiffPool LIN $0.4\pm 1.1$ $1.1\pm 1.9$ $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $22.2\pm 14.3$ $18.9\pm 15.9$ $0.0\pm 0.1$ $2.4\pm 2.1$ OOM
MLP $9.0\pm 1.2$ $6.7\pm 1.0$ $18.9\pm 1.4$ $1.6\pm 0.4$ $3.1\pm 0.9$ $64.5\pm 4.8$ $49.1\pm 4.4$ $9.5\pm 1.2$ $25.8\pm 2.6$ OOM
GCN $39.8\pm 2.0$ $18.2\pm 1.3$ $\mathbf{23.7\pm 1.5}$ $37.0\pm 8.0$ $53.8\pm 7.5$ $73.3\pm 2.0$ $\mathbf{62.4\pm 3.6}$ $33.5\pm 1.2$ $35.7\pm 3.8$ OOM
GIN $30.9\pm 3.6$ $21.3\pm 2.3$ $16.3\pm 3.9$ $14.5\pm 3.2$ $32.8\pm 4.6$ $50.3\pm 3.9$ $41.0\pm 3.9$ $31.6\pm 2.8$ $11.3\pm 3.1$ OOM
SAGE $38.3\pm 2.4$ $16.4\pm 1.4$ $21.9\pm 1.5$ $3.1\pm 0.3$ $16.7\pm 6.1$ $71.4\pm 3.4$ $\underline{59.2\pm 3.3}$ $29.6\pm 2.2$ $38.1\pm 2.8$ OOM
MinCut LIN $26.1\pm 2.5$ $16.5\pm 1.3$ $2.2\pm 1.0$ $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $57.5\pm 0.6$ $36.8\pm 0.9$ $30.4\pm 1.2$ $16.0\pm 1.4$ $3.8\pm 3.5$
MLP $34.7\pm 6.6$ $20.5\pm 0.9$ $5.1\pm 2.4$ $0.0\pm 0$ 0.0 $5.2\pm 9.9$ $53.9\pm 6.4$ $29.4\pm 9.9$ $39.6\pm 1.3$ $28.1\pm 4.2$ $8.0\pm 2.7$
GCN $26.2\pm 1.5$ $12.0\pm 0.7$ $11.0\pm 2.8$ $15.9\pm 21.2$ $45.2\pm 13.6$ $44.7\pm 3.3$ $32.2\pm 6.8$ $26.2\pm 0.7$ $32.1\pm 0.5$ $33.9\pm 4.7$
GIN $39.3\pm 3.8$ $18.9\pm 0.7$ $19.8\pm 4.3$ $6.8\pm 6.8$ $35.5\pm 19.3$ $32.7\pm 19.5$ $51.7\pm 5.4$ $6.9\pm 4.5$ $1.3\pm 6.2$ $10.3\pm 7.3$
SAGE $31.5\pm 1.5$ $16.2\pm 0.8$ $15.5\pm 0.3$ $41.6\pm 0.6$ $46.9\pm 0.5$ $47.1\pm 0.5$ $30.5\pm 2.4$ $35.9\pm 0.7$ $32.3\pm 0.5$ $33.0\pm 1.5$
Ortho LIN $7.0\pm 0.7$ $5.6\pm 0.7$ $16.5\pm 0.4$ $18.8\pm 9.0$ $24.7\pm 11.4$ $45.4\pm 2.0$ $26.7\pm 1.5$ $15.6\pm 1.1$ $29.6\pm 0.9$ $19.1\pm 0.5$
MLP $5.3\pm 0.5$ $3.8\pm 0.4$ $10.6\pm 0.4$ $19.8\pm 5.5$ $30.6\pm 4.0$ $56.3\pm 1.7$ $32.0\pm 1.1$ $5.3\pm 0.5$ $28.8\pm 0.9$ $18.5\pm 0.4$
GCN $22.9\pm 0.9$ $10.9\pm 0.7$ $13.8\pm 0.5$ $40.7\pm 0.9$ $46.0\pm 1.4$ $51.6\pm 1.0$ $32.4\pm 1.0$ $19.2\pm 1.1$ $32.1\pm 0.5$ $34.3\pm 0.7$
GIN $30.7\pm 1.3$ $17.8\pm 0.9$ $13.6\pm 0.5$ $34.0\pm 0.9$ $42.3\pm 0.7$ $44.1\pm 0.6$ $25.0\pm 0.7$ $28.7\pm 1.1$ $29.4\pm 0.6$ $24.1\pm 0.8$
SAGE $18.0\pm 0.9$ $8.4\pm 0.8$ $9.9\pm 0.4$ $35.4\pm 3.9$ $46.3\pm 1.6$ $52.8\pm 1.2$ $34.4\pm 1.1$ $16.3\pm 1.1$ $31.7\pm 0.5$ $24.9\pm 0.3$
Infomap $35.2\pm 0.2$ $23.6\pm 0.1$ $16.0\pm 0.1$ $49.5\pm 0.5$ $57.3\pm 1.1$ $40.6\pm 0.2$ $26.7\pm 0.1$ $35.7\pm 0.3$ $37.9\pm 0.2$ $35.9\pm 0.1$
C.3 Significance of best Neuromap results vs. best baseline
Table 3: Independent two-sample t-test between 25 samples of AMI values for Neuromap vs. the best baseline for $4\sqrt{n}$ hidden features and $s=\sqrt{n}$. The p-values indicate for which datasets the null hypothesis “the samples have the same mean” can be rejected. Blue p-values highlight cases where Neuromap performs significantly better than the best baseline; red p-values highlight cases where the best baseline performs significantly better than Neuromap.
Dataset Best Neuromap Best Baseline p
Cora GIN $46.1\pm 2.0$ NOCD MLP $46.1\pm 1.1$ $0.92$
CiteSeer GIN $23.1\pm 1.2$ NOCD MLP $34.8\pm 0.9$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{3.3e^{-35}}}$
PubMed GIN $24.6\pm 1.7$ DiffPool GCN $23.7\pm 1.5$ $0.05$
PC GCN $52.1\pm 2.3$ NOCD GCN $51.9\pm 1.9$ $0.69$
Photo GCN $65.0\pm 3.8$ NOCD SAGE $63.7\pm 1.6$ $0.14$
CS MLP $78.1\pm 2.0$ NOCD MLP $73.9\pm 1.7$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{4.4e^{-10}}}$
Physics LIN $56.5\pm 1.2$ DiffPool GCN $62.4\pm 3.6$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{1.5e^{-08}}}$
Cora ML SAGE $39.1\pm 2.1$ NOCD MLP $46.5\pm 1.4$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{5.7e^{-18}}}$
Wiki CS SAGE $44.8\pm 2.2$ NOCD GCN $44.2\pm 0.6$ $0.23$
ogb-arxiv SAGE $43.6\pm 1.0$ NOCD GCN $43.8\pm 0.4$ $0.38$
C.4 Significance of Neuromap results vs. Infomap
Table 4: Independent two-sample t-test between 25 samples of AMI values for Neuromap with different (G)NN architectures vs. Infomap for $4\sqrt{n}$ hidden features and $s=\sqrt{n}$. The p-values indicate when the null hypothesis “the samples have the same mean” can be rejected. Blue p-values highlight cases where Neuromap performs significantly better than Infomap; red p-values highlight cases where Infomap performs significantly better than Neuromap.
Dataset LIN MLP GCN GIN SAGE
Cora $0.05$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{6.8e^{-3}}}$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{2.4e^{-8}}}$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{8.1e^{-20}}}$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{9.4e^{-15}}}$
CiteSeer ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{7.5e^{-12}}}$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{4.2e^{-22}}}$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{2.8e^{-20}}}$ $0.06$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{2.1e^{-17}}}$
PubMed $0.65$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{2.1e^{-16}}}$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{4.2e^{-16}}}$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{2.5e^{-18}}}$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{2.2e^{-17}}}$
PC ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{1.8e^{-48}}}$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{4.7e^{-14}}}$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{6.0e^{-6}}}$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{2.4e^{-14}}}$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{3.9e^{-6}}}$
Photo ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{1.8e^{-42}}}$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{6.6e^{-8}}}$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{2.3e^{-10}}}$ $0.48$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{5.2e^{-8}}}$
CS ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{7.6e^{-36}}}$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{1.1e^{-32}}}$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{2.9e^{-32}}}$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{7.5e^{-26}}}$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{5.1e^{-28}}}$
Physics ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{2.7e^{-35}}}$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{1.2e^{-28}}}$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{5.8e^{-31}}}$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{3.2e^{-25}}}$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{2.7e^{-27}}}$
Cora ML ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{9.6e^{-5}}}$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{1.4e^{-3}}}$ $0.01$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{7.0e^{-5}}}$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{1.7e^{-8}}}$
Wiki CS ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{7.0e^{-9}}}$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{8.3e^{-12}}}$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{4.5e^{-5}}}$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{3.9e^{-3}}}$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{3.4e^{-14}}}$
ogb-arxiv ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{1.8e^{-16}}}$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{1.3e^{-11}}}$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{5.4e^{-16}}}$ ${\color[rgb]{0.8515625,0.328125,0.30078125}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.328125,0.30078125}\mathbf{4.6e^{-13}}}$ ${\color[rgb]{0.21484375,0.46875,0.75}\definecolor[named]{pgfstrokecolor}{rgb}{0.21484375,0.46875,0.75}\mathbf{8.9e^{-23}}}$
C.5 Tabulated number of detected communities on real-world networks